pooja195 (pooja195):

I have no idea what to do here. "Is it possible that the car could have accelerated to 55mph within 268 meters if the car can only accelerate from 0 to 60 mph in 15 seconds?"

pooja195 (pooja195):

@nincompoop ;-;

pooja195 (pooja195):

:(

OpenStudy (sachintha):

You need to convert \(mph\) to \(ms^{^-1}\). \(1~mile\approx 1609.34~m\) So, \(Speed = \Large\frac{60\times1609.34}{3600}\normalsize=26.8~ms^{-1}\) You can use \(s=\Large(\frac{v+u}{2})\normalsize t\) to get the distance it car travelled while accelerating from 0 to 60 mph. \(s=\Large(\frac{26.8+0}{2})\normalsize 15\approx 201~m\) So it's possible to accelerate to 55mph within 268 m.

pooja195 (pooja195):

#confused